2. (i) Prove that if n is a positive integer, then n is even if and only if 7n + 4 is even.
(ii) Prove that if n is a positive integer, then n is odd if and only if 5n + 6 is odd.
(iii) Prove that m2 = n2 if and only if m = n or m = -n.
(iv) Prove or disprove that if m and n are integers such that mn = 1, then either m = 1 or else m = -1 and n = -1.
1
Expert's answer
2020-06-22T17:36:33-0400

2.(i)"=>" We prove first the direct implication. Assume "n" is even. Then "n=2k" for some integer "k." Then "7n+4=7(2k)+4=14k+4=2(7k+2)," which is even.


"<=" For the converse, which is “if "7n+4" is even, then "n" is even”, we use a proof by contrapositive.

The contrapositive is: “if "n" is not even (that is, odd), then "7n+4" is not even (that is, odd)”. If "n" is odd, then "n=2k+1," for some integer "k."

Then "7n+4=7(2k+1)+4=14k+7+4=14k+11," which is odd.

Hence we have proven that "n" is even if and only if "7n+4" is even. 


(ii) "=>" We will use a direct proof on “If "n" is odd, then "5n+6" is odd”. Assume "n" is odd, so "n=2k+1" for some integer "k."

Then "5n+6=5(2k+1)+6=10k+5+6" "=10k+11=2(5k+5)+1."

Thus, "5n+6" is odd.


"<=" We now must prove the converse, “If "5n+6" is odd, then "n" is odd.” For this, we will use proof by contrapositive. So the statement becomes “If "n" is not odd, then "5n+6" is not odd.” Assume that "n" is not odd, so "n=2k" for some integer "k."

Then "5n+6=5(2k)+6=10k+6=2(5k+3)."

Thus, "5n+6" is not odd.

Hence we have proven that "n" is odd if and only if "5n+6" is odd. 


(iii) "=>" If "m=n," then "m^2=n^2" follows using basic principles of Algebra.

Similarly, if "m=-n," then "m^2=(-n)^2=n^2" follows using basic principles of Algebra.

"<=" Suppose that "m^2=n^2." Notice that the square of any real number is non-negative, so we can take the square root of each side of this equation in order to obtain: "\sqrt{m^2}=\sqrt{n^2}," or "|m|=|n|." Thus, using basic principles of algebra, either "m=n" or "m=-n."


(iv) "mn=1=>m\not=0, n\not=0."

Case 1: Suppose "|m|>1." Regardless of what "n" is, "mn=1," so "n=\dfrac{1}{m}." But we said that both "m" and "n" are integers. If "|m|>1," then "n" here cannot be an integer. So, contradiction.

Case 2: Suppose "|n|>1." Regardless of what "m" is, "mn=1," so "n=\dfrac{1}{m}." But we said that both "m" and "n" are integers. If "|n|>1," then "m" here cannot be an integer. So, contradiction.

Case 3: Suppose "m=1" and "n=-1." Clearly then "mn=-1." So, contradiction.

Case 4: Suppose "m=-1" and "n=1." Clearly then "mn=-1." So, contradiction.

Case 5: Suppose "m=1" and "n=1." Then "mn=1."

Case 6: Suppose "m=-1" and "n=-1." Then "mn=1."

What techniques did we use? Cases and contradiction